OpenStudy (anonymous):

Points A and B are the foci of a conic section. Two points on the conic section, C and D, have distances to the foci A and B as shown. CA=5 CB=10 DA=8 In order for the points C and D to be on the same ellipse, the distance would have to equal ?

OpenStudy (tkhunny):

Ellipse - Sum of the two distances from the foci is constant. d(CA) + d(CB) = 5 + 10 = 15 d(DA) + d(DB) = 8 + ?? = 15

OpenStudy (anonymous):

Thank you!